Someone please show me how to do these so I can help my son
-1/5y=-6
3/4z= -5 1/2
-2/3t=72

Answer by Alan3354(69443)

-1/5y=-6
Assuming you mean (-1/5)y and not -1/(5y):
Multiply by -1
(1/5)y = 6
Multiply by 5
y = 30
-----------
3/4z= -5 1/2
(3/4)z = -11/2
Multiply by 4
3z = -22
Divide by 3
z = -22/3
------------
-2/3t=72
Multiply by 3
-2t = 216
Divide by -2
t = -108